the height of the water H in feet at a boat dock t hours after 6 a.m. is given by H=10+sin pi/4 t. what is the height of the water at 10 a.m. ?
a. 5 feet
b. 12 feet
c. 15 feet
d. 10 feet

Found 2 solutions by Alan3354, rothauserc:
Answer by Alan3354(69443)

the height of the water H in feet at a boat dock t hours after 6 a.m. is given by H=10+sin pi/4 t. what is the height of the water at 10 a.m. ?
--------------
t = 10-6 = 4
h = 10 + sin(pi*t/4) = 10 + sin(pi)
h = 10 feet

Answer by rothauserc(4718)

H = 10 + sin((pi/4)*t)
t = 10 - 6 = 4
H = 10 + sin((pi/4)*4)
H = 10 + sin(45*4)
H = 10 + 0 = 10
answer is d.